the same number of each coin have a total of $6.00. there's nickels , dimes and quarters. how many of each coin are there
Answer by richwmiller(17219)

25x+5x+10x=600
40x=600
x=15 of each
check
25*15+10*15+5*15=600
375+150+75=600
ok